# Collapsing in polygonal dynamics

### v2

@@ 6.1 Definition and results
Now we can compute the infinitesimal monodromy and the coordinate of the collapse points [^13] The polynomial for the collapse points was first found experimentally by Paul Melotti, when working on discrete holomorphic functions.. Figure 4 depicts the convergence of one randomly chosen closed $5$-gon, under the iteration of some non-trivial element of $\Gamma_{5}$, towards one of the collapse points, as predicted by conjecture 3.2.
@@
$M^{\prime}=\sum_{i=1}^{n}\eta\mu_{i}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
@@
So, up to an unimportant sign, we get back the same formula as in proposition 5.6. This could be expectable, since they both arise from the theory of discrete holomorphic functions.
@@
$\frac{c_{i}^{\prime}(1)}{c_{i}}=\frac{c_{i}\eta\mu_{i}}{c_{i}}=\eta\mu_{i},$
@@
Figure 4: Python simulation of the evolution of one randomly chosen closed $5$-gon in $\mathbb{P}^{1}(\mathbb{C})$ with randomly chosen discrete curvature, under the iteration of the sequence of flips $\phi_{5}\circ\dots\circ\phi_{1}$ (applied 25 times). The initial vertices are represented with red crosses. A new edge is drawn each time a flip is done, and the colour indicates time (from blue to red). There is a rapid spiraling towards one of the collapse points determined by corollary 3.4.
@@ 6.2 Special configurations
Here, we investigate some specific cases of staircase cross-ratio dynamics for which we can write explicitly the dynamic. Their motion on the moduli space is periodic, meaning that it amounts on the initial space to iterate a projective transformation (of any kind). They can be seen as a special case of discrete holomorphic functions, depicted in figure 5. The idea comes from [31], and it holds on any algebraicly closed field of characteristic 0. For an example of generic orbit in $\mathbb{P}^{1}(\mathbb{C)}$, see the simulation in figure 4.
First, let’s see that how a parabolic transformation can appear. Let’s consider a closed $n$-gon
@@
Let’s go back to conjecture 3.2. In the case of periodic motion of the moduli space, any type of Möbius transformation can appear.
@@
    If $q$ is a root of unity, then the dynamic is periodic (finite order elliptic transformation);
@@
In all of this cases, we can deform the polygons via the scaling symmetry to get families of twisted $n$-gons. The generic case is the first, which is coherent with conjectures 3.2, but the other ones remind us that it won’t hold everywhere.
@@
This work was partially funded by the French National Research Agency under the project DynAtrois (ANR-24-CE40-1163)

### v3

@@ 6.1 Definition and results
Now we can compute the infinitesimal monodromy and the coordinate of the collapse points [^19] The polynomial for the collapse points was first found experimentally by Paul Melotti when working on discrete holomorphic functions.. Figure 6 depicts the convergence of a random closed $5$-gon, under the iteration of some non-trivial element of $\Gamma_{5}$, towards one of the collapse points, as predicted by conjecture 3.4.
@@
$M^{\prime}(1)=\sum_{i=1}^{n}\eta\mu_{i}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
@@
We get back the same formula as in proposition 5.6. This was expectable, since they both arise from the theory of discrete holomorphic functions.
@@
$\frac{c_{i}^{\prime}(1)}{c_{i}(1)}=\frac{c_{i}\eta\mu_{i}}{c_{i}}=\eta\mu_{i},$
@@
Figure 6: Python simulation of the evolution of a random closed $5$-gon in $\mathbb{P}^{1}(\mathbb{C})$ with random discrete curvature, under the iteration of the sequence of flips $\phi_{5}\circ\dots\circ\phi_{1}$ (applied 25 times). The initial vertices are represented with red crosses. A new edge is drawn each time a flip is done, and the colour indicates time (from blue to red). There is a rapid spiraling towards one of the collapse points determined by theorem 3.5.
@@ 6.2 Special configurations
Here, we investigate some specific cases of staircase cross-ratio dynamics for which we can write explicitly the dynamic. Their motion on the moduli space is periodic, meaning that it amounts on the initial space to iterate a projective transformation (of any kind). They can be seen as a special case of discrete holomorphic functions, depicted on figure 7. The idea comes from [leitenbergerIteratedHarmonicDivisions2016a], and it holds on any algebraicly closed field of characteristic 0. For an example of generic orbit in $\mathbb{P}^{1}(\mathbb{C)}$, see the simulation on figure 6.
First, let’s see that how a parabolic transformation can appear. Let’s consider the closed $n$-gon
@@
Let’s go back to theorem 3.1. In the case of periodic motion of the moduli space, any type of Möbius transformation can appear.
@@
    if $q\neq 1$ is a root of unity, then the dynamic is periodic (finite order elliptic transformation);
@@
In all of this cases, we can deform the polygons via the scaling symmetry to get families of twisted $n$-gons whose vertices also collapse.
@@
This work was partially funded by the French National Research Agency under the project DynAtrois (ANR-24-CE40-1163).